An ice hockey rink is a rectangle with a perimeter of 174 meters. The length is 9 meters more than twice the width. Find the length and width...
this problem is frustrating...I need it in steps please :-) I could really use some help fellas! :-| eeck
thanks so much!!!

Found 2 solutions by longjonsilver, queenofit:
Answer by longjonsilver(2297)

Let width be x metres

so, length is "9 metres more than twice the width" --> 9+2x

So, perimeter is x + x + (9+2x) + (9+2x) = 174
6x + 18 = 174
6x = 156
x = 156/6
x = 26m

width = 26m
length is therefore 9+2(26) --> 9+52 --> 61m

jon.

Answer by queenofit(120)

Fellas?????women like math as well, I should let you drown in this math problem, but I won't, just kidding really, just thought you needed a laugh, you sounded so serious....
For this, and other problems like this you need to know several things before you can start. you are looking for the perimeter, so you must know the formula for the perimeter of a rectangle. that is 2L+2W=174meters in this case. You are looking for the L and the W. So, to solve, the easiest way is to look for the variable that does not have a number involved with it. In this case that is your W. So we will solve for W. so this is what you know. always write this down first when you solve a problem like this.
2L+2W=174
W=W (this is a single side)
L= 9+2w (this is a single side)
put this in your above equation. Notice that when you put it into the equation you have to multiply both your lenth and width by 2 to satisfy your equation for perimeter. The W and L above is only for one side, but in the equation you must have both sides, that is why you mult. by 2 in the equation. I did it below.
2(9+2w)+2w=174 (distribute your multiplication next)
18+4w + 2w = 174 (combine like terms next)
18+ 6w = 174 ( move your 18 to the left using properties of algebra)
6w=174-18
6w=156 (divide by 6 to get w)
w= 26
To get the length your equation said it was 9 more than twice the width
9+2(26)=61
to check
2w+2L=174
2(26)+2(61)=174
52+122=174
Hope this helps, try not to stress so much, it will come in time, it takes practice.